# UniAudio: An Audio Foundation Model Toward Universal Audio Generation

### v5

@@ B.3 Singing Voice Synthesis
Diffsinger (Liu et al., 2022) 3.94$\pm$0.02 4.05$\pm$0.06
Make-A-Voice (Huang et al., 2023b) 3.96$\pm$0.03 4.04$\pm$0.05
@@ B.4 Text-to-sound and text-to-music generation
The text-to-sound generation task has attracted great interest in audio research. Following Diffsound (Yang et al., 2023c), most of the methods evaluate their systems on the AudioCaps (Kim et al., 2019) test set. However, we found that if the training data includes the AudioCaps data, the model is easy to overfit with AudioCaps. As a result, the best performance can be obtained when the model only trains on the Audiocaps. In this study, we conduct a zero-shot evaluation on the Cloth test set (Drossos et al., 2020). Table 13 shows the results. We can see that UniAudio obtains better performance than Diffsound and AudioLDM. Compared to recent SOTA models, such as Tango and Make-an-Audio 2, UniAudio also gets comparable performance. For the text-to-music task, we follow MusicGen (Copet et al., 2023), evaluating our methods on MusicCaps (Agostinelli et al., 2023). Compared with previous SOTAs, UniAudio gets a comparable performance with other models. From the MOS evaluation performance, we can see that MusicGen is better than our current models. We speculate one of the reasons is that MusicGen uses a large-scale high-quality dataset (20k hours).
@@ B.5 Audio Edit
Audio edit aims to edit the original audio based on Human’s instruction. AUDIT (Wang et al., 2023d) is the SOTA model in audio edit task, which designs a data simulation strategy to get triplet training and test data (e.g., {audio, audio, text}). The authors set 5 different tasks, including adding, dropping, replacing, inpainting and super-resolution, and simulated large-scale data for each task. To validate that our pre-trained model can be fine-tuned with small-scale data, we choose adding, dropping and super-resolution tasks to fine-tune simultaneously. To finish the fine-tuning process, we define a new task label: Audit_task. The experimental results as Table 14 shows. We can observe that: (1) UniAudio can get better performance with the previous SOTA model. (2) Fine-tuning pre-trained UniAudio can get better performance than training it from scratch, which further validates the effectiveness of pre-training a model on large-scale training data.
@@ B.6 Instructed TTS
Using instruction to guide speech synthesis has received great attention (Guo et al., 2023; Yang et al., 2023a). In this part, we fine-tune the UniAudio model on the PromptSpeech (Guo et al., 2023) dataset. Furthermore, we also try to train a UniAudio model from scratch with the PromptSpeech dataset. Different from previous works that designed special style encoders to capture the style information from text descriptions, we directly use the T5 text encoder to extract representations from text and then combine it with the phoneme sequence input to the UniAudio, which is more convenient. [^16] Note that the authors of PromptTTS (Guo et al., 2023) told us their objective metrics tools, checkpoints, and generated samples have been lost due to the machine errors. Thus we cannot fairly compare with them. Table 15 shows the results, we can see that UniAudio has good performance in terms of style control and speech quality when compared with the ground truth samples.
@@ C.1 The influence of multi-task training
Text-to-Speech UniAudio (Single) SIM()($\uparrow$)( ↑ ) / WER()($\downarrow$)( ↓ ) 0.64 / 2.4 MOS()($\uparrow$)( ↑ ) / SMOS(($\uparrow$( ↑) 3.77$\pm$0.06 / 3.46$\pm$0.10
@@
Voice Conversion UniAudio (Single) SIM()($\uparrow$)( ↑ ) / WER()($\downarrow$)( ↓ ) 0.84 / 5.4 MOS()($\uparrow$)( ↑ ) / SMOS(($\uparrow$( ↑) 3.45$\pm$0.07 / 3.44$\pm$0.07
@@
Speech Enhancement UniAudio (Single) PESQ()($\uparrow$)( ↑ ) / VISQOL()($\uparrow$)( ↑ ) / DNSMOS()($\uparrow$)( ↑ ) 2.35 / 2.30 / 3.45 MOS(($\uparrow$( ↑) 3.65$\pm$0.08
@@
Target Speaker Extraction UniAudio (Single) PESQ()($\uparrow$)( ↑ ) / VISQOL()($\uparrow$)( ↑ ) / DNSMOS()($\uparrow$)( ↑ ) 1.97 / 1.61 / 3.93 MOS(($\uparrow$( ↑) 3.58$\pm$0.08
@@
Singing Voice Synthesis UniAudio (Single) - - MOS(($\uparrow$( ↑) / SMOS(($\uparrow$( ↑) 4.14$\pm$0.07 / 4.02$\pm$0.02
@@
Text-to-Sound UniAudio (Single) FAD ()($\downarrow$)( ↓ ) / KL ()($\downarrow$)( ↓ ) 3.84 / 2.7 OVL ()($\uparrow$)( ↑ ) / REL ()($\uparrow$)( ↑ ) 60.0$\pm$2.1 / 61.2$\pm$1.8
@@
Text-to-Music UniAudio (Single) FAD ()($\downarrow$)( ↓ ) / KL ()($\downarrow$)( ↓ ) 5.24 / 1.8 OVL ()($\uparrow$)( ↑ ) / REL ()($\uparrow$)( ↑ ) 64.4$\pm$2.1 / 66.2$\pm$2.4
@@
Audio Edit UniAudio (single) FD ()($\downarrow$)( ↓ ) / KL ()($\downarrow$)( ↓ ) 19.82 / 0.92 - -
@@
Speech Dereverb. UniAudio (single) PESQ()($\uparrow$)( ↑ ) / DNSMOS()($\uparrow$)( ↑ ) 1.23 / 3.18 - -
@@
Instructed TTS UniAudio (single) - - MOS()($\uparrow$)( ↑ ) / SMOS()($\uparrow$)( ↑ ) 3.62$\pm$0.07 / 3.67$\pm$0.08
@@
Speech Edit UniAudio (single) MCD ()($\downarrow$)( ↓ ) 5.26 MOS()($\uparrow$)( ↑ ) 3.73$\pm$0.07
@@ Appendix D Why UniAudio Can Work Well?
(1) Deterministic latent space: we formulate different modalities into a deterministic latent space (fixed vocabulary) by tokenization. Different tokens can be seen as specific ’words’, and we can use a next-token prediction strategy to train the model. Similar to GPT-series (Radford et al., 2018; 2019), such strategy creates the opportunity for the model to learn the intrinsic properties of audio and the interrelationship between audio and other modalities.
@@
We can see that the difference in task formulation for TTS and VC is that they use different ways to denote the phonetic information. In essence, they carry the same phonetic information. The difference is that semantic tokens include the duration information. Thus we can view the phoneme sequence as a special semantic sequence that drops the duration information. Such dropping operation is widely used as a data augmentation strategy (Park et al., 2019).
@@ Appendix E The details of Audio Codec Models
(Veaux et al., 2017) (Drossos et al., 2020) (Agostinelli et al., 2023) (Zhang et al., 2022) -
@@
In this part, we give more details about our neural audio codec model in Section 2.1.1. We adopt a similar encoder-decoder framework with the Encodec model, the difference includes: (1) we replace the multi-scale STFT-based (MS-STFT) discriminator as our multi-scale Mel-based discriminator. (2) We rewrite the vector quantization implementation  [^17] Please refer to our source code to find the details. based on Encodec’s open-source version  [^18] https://github.com/facebookresearch/encodec/blob/main/encodec/quantization/core_vq.py, making it more suitable for DDP training. Figure 5 shows the details of the mel-based discriminator. We combine the mel-spectrogram and log-mel-spectrogram features and then input them into a network consisting of several convolutional layers. Our motivation is that the mel-spectrogram has a strong intrinsic inductive bias, especially for sounds and music-related audio (the SOTA sounds or music classification systems are based on the log-mel-spectrogram in the literature.). Thus, we speculate that choosing a mel-spectrogram-based discriminator can better promote high-fidelity audio reconstruction. In our experiments, we use 6 different discriminators with different configurations  [^19] In our experiments, we find the mel-based discriminator brings better reconstruction performance when we train a universal neural audio codec.. Specifically, we set the hidden_dim as {64, 128, 256, 512, 512, 512} and the hop length as {32, 64, 128, 256, 512, 1024}. We train the neural audio codec model based on the Librilight and AudioSet datasets. Table 18 demonstrates that the neural codec model adopted in this work outperforms prior Encodec (Défossez et al., 2022).
@@ Appendix F Subjective Evaluation
For sound and music generation tasks, we follow AudioGen (Kreuk et al., 2022) and MusicGen (Copet et al., 2023) to evaluate (1) overall quality (OVL), and (2) relevance to the text input (REL).

### v6

@@ B.3 Singing Voice Synthesis
Diffsinger (liu2022diffsinger) 3.94$\pm$0.02 4.05$\pm$0.06
Make-A-Voice (makeavoice) 3.96$\pm$0.03 4.04$\pm$0.05
@@ B.4 Text-to-sound and text-to-music generation
The text-to-sound generation task has attracted great interest in audio research. Following Diffsound (diffsound), most of the methods evaluate their systems on the AudioCaps (audiocaps) test set. However, we found that if the training data includes the AudioCaps data, the model is easy to overfit with AudioCaps. As a result, the best performance can be obtained when the model only trains on the Audiocaps. In this study, we conduct a zero-shot evaluation on the Cloth test set (drossos2020clotho). Table 11 shows the results. We can see that UniAudio obtains better performance than Diffsound and AudioLDM. Compared to recent SOTA models, such as Tango and Make-an-Audio 2, UniAudio also gets comparable performance. For the text-to-music task, we follow MusicGen (musicgen), evaluating our methods on MusicCaps (musiclm). Compared with previous SOTAs, UniAudio gets a comparable performance with other models. From the MOS evaluation performance, we can see that MusicGen is better than our current models. We speculate one of the reasons is that MusicGen uses a large-scale high-quality dataset (20k hours).
@@ B.5 Audio Edit
Audio edit aims to edit the original audio based on Human’s instruction. AUDIT (wang2023audit) is the SOTA model in audio edit task, which designs a data simulation strategy to get triplet training and test data (e.g., {audio, audio, text}). The authors set 5 different tasks, including adding, dropping, replacing, inpainting and super-resolution, and simulated large-scale data for each task. To validate that our pre-trained model can be fine-tuned with small-scale data, we choose adding, dropping and super-resolution tasks to fine-tune simultaneously. To finish the fine-tuning process, we define a new task label: Audit_task. The experimental results as Table 12 shows. We can observe that: (1) UniAudio can get better performance with the previous SOTA model. (2) Fine-tuning pre-trained UniAudio can get better performance than training it from scratch, which further validates the effectiveness of pre-training a model on large-scale training data.
@@ B.6 Instructed TTS
Using instruction to guide speech synthesis has received great attention (guo2023prompttts; yang2023instructtts). In this part, we fine-tune the UniAudio model on the PromptSpeech (guo2023prompttts) dataset. Furthermore, we also try to train a UniAudio model from scratch with the PromptSpeech dataset. Different from previous works that designed special style encoders to capture the style information from text descriptions, we directly use the T5 text encoder to extract representations from text and then combine it with the phoneme sequence input to the UniAudio, which is more convenient. [^7] Note that the authors of PromptTTS (guo2023prompttts) told us their objective metrics tools, checkpoints, and generated samples have been lost due to the machine errors. Thus we cannot fairly compare with them. Table 13 shows the results, we can see that UniAudio has good performance in terms of style control and speech quality when compared with the ground truth samples.
@@ C.1 The influence of multi-task training
[…]
/ SMOS(($\uparrow$( ↑) 3.77$\pm$0.06 / 3.46$\pm$0.10
@@
[…]
/ SMOS(($\uparrow$( ↑) 3.45$\pm$0.07 / 3.44$\pm$0.07
@@
[…]
/ VISQOL()($\uparrow$)( ↑ ) / DNSMOS()($\uparrow$)( ↑ ) 2.35 / 2.30 / 3.45 MOS(($\uparrow$( ↑) 3.65$\pm$0.08
@@
[…]
/ VISQOL()($\uparrow$)( ↑ ) / DNSMOS()($\uparrow$)( ↑ ) 1.97 / 1.61 / 3.93 MOS(($\uparrow$( ↑) 3.58$\pm$0.08
@@
[…]
/ SMOS(($\uparrow$( ↑) 4.14$\pm$0.07 / 4.02$\pm$0.02
@@
[…]
/ REL ()($\uparrow$)( ↑ ) 60.0$\pm$2.1 / 61.2$\pm$1.8
@@
[…]
/ REL ()($\uparrow$)( ↑ ) 64.4$\pm$2.1 / 66.2$\pm$2.4
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ Appendix D Why UniAudio Can Work Well?
(1) Deterministic latent space: we formulate different modalities into a deterministic latent space (fixed vocabulary) by tokenization. Different tokens can be seen as specific ’words’, and we can use a next-token prediction strategy to train the model. Similar to GPT-series (radford2018improving; radford2019language), such strategy creates the opportunity for the model to learn the intrinsic properties of audio and the interrelationship between audio and other modalities.
@@
We can see that the difference in task formulation for TTS and VC is that they use different ways to denote the phonetic information. In essence, they carry the same phonetic information. The difference is that semantic tokens include the duration information. Thus we can view the phoneme sequence as a special semantic sequence that drops the duration information. Such dropping operation is widely used as a data augmentation strategy (specaugment).
@@ Appendix E The details of Audio Codec Models
(vctk) (drossos2020clotho) (musiclm) (m4singer) -
@@
In this part, we give more details about our neural audio codec model in Section LABEL:sec:audio. We adopt a similar encoder-decoder framework with the Encodec model, the difference includes: (1) we replace the multi-scale STFT-based (MS-STFT) discriminator as our multi-scale Mel-based discriminator. (2) We rewrite the vector quantization implementation  [^8] Please refer to our source code to find the details. based on Encodec’s open-source version  [^9] https://github.com/facebookresearch/encodec/blob/main/encodec/quantization/core_vq.py, making it more suitable for DDP training. Figure 4 shows the details of the mel-based discriminator. We combine the mel-spectrogram and log-mel-spectrogram features and then input them into a network consisting of several convolutional layers. Our motivation is that the mel-spectrogram has a strong intrinsic inductive bias, especially for sounds and music-related audio (the SOTA sounds or music classification systems are based on the log-mel-spectrogram in the literature.). Thus, we speculate that choosing a mel-spectrogram-based discriminator can better promote high-fidelity audio reconstruction. In our experiments, we use 6 different discriminators with different configurations  [^10] In our experiments, we find the mel-based discriminator brings better reconstruction performance when we train a universal neural audio codec.. Specifically, we set the hidden_dim as {64, 128, 256, 512, 512, 512} and the hop length as {32, 64, 128, 256, 512, 1024}. We train the neural audio codec model based on the Librilight and AudioSet datasets. Table 16 demonstrates that the neural codec model adopted in this work outperforms prior Encodec (encodec).
@@ Appendix F Subjective Evaluation
For sound and music generation tasks, we follow AudioGen (kreuk2022audiogen) and MusicGen (musicgen) to evaluate (1) overall quality (OVL), and (2) relevance to the text input (REL).